OpenStudy (anonymous):

use roots to rewrite and simplify 641/2

OpenStudy (mathstudent55):

You mean \(\LARGE 64^{\frac{1}{2} }\) ?

OpenStudy (anonymous):

64 and 1/2

OpenStudy (mathstudent55):

Remember that \(a ^ {\frac{1}{n}} = \sqrt[n] {a} \)

OpenStudy (mathstudent55):

You mean 64 to the 1/2 power?

OpenStudy (anonymous):

uhm further step please

OpenStudy (anonymous):

yea 64 to the 1/2

OpenStudy (mathstudent55):

From the formula above, \(64^{\frac{1}{2}} = \sqrt{64} = 8\)